2x+2y-8z=8
5x+y-z=8
-x-y+4z=2

Answer by FrankM(1040)

2x+2y-8z=8
5x+y-z=8
-x-y+4z=2
This is 3 equations in 3 unknowns, i.e. 3 variables. We need to manipulate to get to 2 equations /2 unknowns, by combining 2 at a time.
2x+2y-8z=8 (EQ1)
-2x-2y+8z=4 (EQ3 multiplied by 2)
All terms cancel, adding to zero. There is no unique solution.